A market research firm supplies manufacturers with estimates of the retail sales of their products from samples of retail stores. Marketing managers are prone to look at the estimate and ignore sampling error. Suppose that an SRS of 60 stores this month shows mean sales of 53 units of a small appliance, with a standard deviation of 12 units. During the same point in time last year, an SRS of 58 stores gave mean sales of 50 units with standard deviation 10 units. An increase from 50 to 53 is a rise of 6%. The marketing manager is happy because sales are up 6%.

Use the two-sample t procedure to give a 95% confidence interval for the difference in mean number of units sold at all retail stores.
Answer #1

SP nl n2- 2

V/ (60-1) × 122 + (58-1) × 102 = 1 1 .0625 6058-2

SE=s_{P}\sqrt{1/n1+1/n2}

11.0625/1/60 +1/58 - 2.0358

\alpha = 0.05

ndf = n1+n2 -2 = 60 + 58 - 2 = 116

From Table, critical values of t = \pm 1.9806

Confidence Interval:

(53 - 50) \pm (1,9806 X 2.0358)

= 3 \pm 4.0321

= ( -1.0321, 7.0321)

Confidence Interval:

-1.0321 < \mu 1-\mu 2 < 7.0321